OpenStudy (anonymous):

Need help with a derivative: \[\frac{ d }{ dt } e ^{kcos(2t)}\] I think this should be : \[ke ^{kcos(2t)} *\frac{ d }{ dt } kcos(2t)\] Which is then : \[ke ^{kcos(2t)} * -ksin(2t)\]

OpenStudy (anonymous):

but apparently this is wrong. Does anyone see where I went wrong?

OpenStudy (vishweshshrimali5):

You did several mistakes in differentiating the given function.

OpenStudy (vishweshshrimali5):

First of all while differentiating the first time you did not follow the chain rule properly.

OpenStudy (anonymous):

oh ya, i should have brough the whole thing down, not just a k....

OpenStudy (vishweshshrimali5):

It should have been: \[\large{\cfrac{ d }{ dt } e ^{kcos(2t)}}\] \[\large{=e^{k\cos(2t)} * \cfrac{d}{dt}(k\cos(2t))}\]

OpenStudy (anonymous):

oh

OpenStudy (vishweshshrimali5):

Now, while calculating derivative of kcos(2t), you did the same mistake. You differentiated cos correctly but forgot to differentiate the angle i.e. (2t).

OpenStudy (vishweshshrimali5):

So, it should have been: \[\large{\cfrac{d}{dt}(k\cos(2t))}\] \[\large{= k*(-\sin(2t))*2}\] \[\large{=-2k\sin(2t)}\]

OpenStudy (vishweshshrimali5):

Did you get it ?

OpenStudy (anonymous):

ya, i didnt realize the chain rule applied to the angle as well...

OpenStudy (vishweshshrimali5):

Its okay.

OpenStudy (anonymous):

thanks for clearing that up, its been quite a while :D

OpenStudy (vishweshshrimali5):

It was my pleasure to be of some help. Best of luck for further questions and good night (according to Indian Standard Time)!!!